OpenStudy (anonymous):

plz help me prove that he set S ={q belongs to Q : q > 0 and q2 < 2}This set has no largest number.

ganeshie8 (ganeshie8):

Do you mean \[S = \{q \in \mathbb{Q} : q \gt 0~ \mathrm{and }~q^2 \lt 2\}\] ?

OpenStudy (zzr0ck3r):

wow you get a medal just for decoding.

ganeshie8 (ganeshie8):

:) this looks same as proving x^2=2 is not solveable in rational numbers ?

OpenStudy (zzr0ck3r):

Can you use density of \(\mathbb{Q}\) in \(\mathbb{R}\)?

OpenStudy (zzr0ck3r):

its an non empty open interval, so density would take care of it...

OpenStudy (ikram002p):

ok as i said before its the same as proving rational number are infinite, i got dull for a second cuz of ganesh comment but now ok :P

OpenStudy (ikram002p):

so basically , show that the set of q^2 is an infinite countable subset of rational

ganeshie8 (ganeshie8):

nvm my comments im an amateur in topology/analysis lol

OpenStudy (zzr0ck3r):

How does this prove it?

OpenStudy (ikram002p):

its prove cuz of this :- infinite=no absolute supremum .

OpenStudy (ikram002p):

ok continue @zzr0ck3r

OpenStudy (zzr0ck3r):

er \(x_0\in (x,\sqrt{2})\) and we are done

OpenStudy (zzr0ck3r):

but just because something is infinite does not mean a proper subset is

OpenStudy (ikram002p):

humm , well i cant see ur way yet :O what i mean is this :- show that q^2 is closed on rational xD

OpenStudy (ikram002p):

this is first step

OpenStudy (zzr0ck3r):

are we reading the same question? I am reading this. Show that there is no largest element in \((0,\sqrt{2})\)

OpenStudy (ikram002p):

:)

OpenStudy (zzr0ck3r):

best way to show that \(\mathbb{Q}\) is countable is \(f(\frac{m}{n}) =(m,n)\in \mathbb{N}\times \mathbb{N}-\{0\}\) where \(\frac{m}{n}\) is in reduced form. I only say this because I just learned it today, and its quick and easy:)

OpenStudy (zzr0ck3r):

p.s. that was bad notation when I put (0,sqrt2) I mean the rationals in there....

ganeshie8 (ganeshie8):

zzr your first argument looks a lot like \(\forall x \in (0, \sqrt{2}), \) \[x\lt \dfrac{x+\sqrt{2}}{2} \lt \sqrt{2} \] (i had to convert like this as im not familiarty with topology terms like dense etc.. )

OpenStudy (zzr0ck3r):

let me make more clear

OpenStudy (perl):

you cant use sqrt(2) in your proof, since sqrt(2) is not defined (yet)

ganeshie8 (ganeshie8):

thats a good point

OpenStudy (perl):

and we are working only with q , such that 0<q^2 < 2 ,

OpenStudy (zzr0ck3r):

consider the set \(S=\{x\in \mathbb{Q}| x\in (0,\sqrt{2})\}\). Suppose by way of contradiction there exists a greatest element \(x\) in \(S\). Then by the density of the rationals there is a \(y\in \mathbb{Q}\) s.t. \(x<y<\sup S\). This is a contradiction.

ganeshie8 (ganeshie8):

maybe lets assume for now that we have created Reals already to continue with these cool proofs :)

OpenStudy (ikram002p):

well i said my view , in case u learned it yet or not simply show that q^2<2 is closed in Q , and then show its infinite for q>0 done .

OpenStudy (zzr0ck3r):

This of course only works if they have proved that the rationals are dense in the reals. i.e. between any two numbers a<b there is a rational number c s.t. a<c<b I wish the user would tell us what they can use:)

OpenStudy (zzr0ck3r):

yeah I don't follow. Are we talking the topological definition of closed?

OpenStudy (ikram002p):

i liked this one clear and proper. http://prntscr.com/5cqseu also arkhamedian use same thing , about rational

OpenStudy (zzr0ck3r):

in his bath?

ganeshie8 (ganeshie8):

still going thru previous replies.. just a quick q\[f(\frac{m}{n}) =(m,n)\in \mathbb{N}\times \mathbb{N}-\{0\}\] this is like cantor's diagonalization proof where they build a matrix and count diagonally right ?

ganeshie8 (ganeshie8):

the product of two countable sets is countable or somehting..

OpenStudy (zzr0ck3r):

yeah. Its similar, but he didn't talk about 1-1 in that proof

OpenStudy (zzr0ck3r):

product and sum

OpenStudy (perl):

you might just want to clarify that your sup S is a rational number

OpenStudy (zzr0ck3r):

why?

OpenStudy (perl):

it might be confused with the square root 2

OpenStudy (zzr0ck3r):

Q is dense in R

OpenStudy (zzr0ck3r):

the supremum need not be in the set, and thus might not be rational.

OpenStudy (zzr0ck3r):

I assume we are talking about R being our universe here....

OpenStudy (perl):

the set S is only rational numbers but the proof hinges on density of rationals, which hasnt been proven, and sqrt(2) which is undefined . so you are taking a lot for granted in your proof

OpenStudy (zzr0ck3r):

are you in the class?

OpenStudy (zzr0ck3r):

should we start from set axioms? I have said about 5 times now if you simply read that this proof is valid if and only if density is allowed.

OpenStudy (perl):

sorry i didn't read the entire thread . (take it easy)

OpenStudy (zzr0ck3r):

I am easy :) lol Just pointing out what I already said

OpenStudy (perl):

but sqrt(2) was not given in the original statement. im not sure if you should treat it like a number

OpenStudy (blacksteel):

I feel like this is really easy to prove by contradiction. Basically, this boils down to proving that there is no greatest element\[0 < q < \sqrt{2}, q \in Q\] So suppose x is the greatest such value. Then let\[0 < y < \sqrt{2} - x, y \in Q\] Then x + y is a value in Q that satisfies the conditions for q.

OpenStudy (zzr0ck3r):

then remove sqrt of two and call is S. its the same set....

OpenStudy (zzr0ck3r):

this is the same thing I did .... you are using density

ganeshie8 (ganeshie8):

yes its same as showing the average of x and sqrt(2) is greater than x

OpenStudy (zzr0ck3r):

if you can show such a y there is no need to do the sqrt(2) -x just pick y between x and sqrt(2)

OpenStudy (zzr0ck3r):

or the sup of the set if we don't like irrationals...

OpenStudy (zzr0ck3r):

but the sup of the set is irrational so I am unclear what the point was there...

ganeshie8 (ganeshie8):

hey i found this proof that uses only rational numbers, but i dont get it completely

OpenStudy (ikram002p):

i remember same page in my book :O lol

ganeshie8 (ganeshie8):

looks it is using complicated algebra instead of density and other sophisticated tools

OpenStudy (zzr0ck3r):

prob cooler.... This is why I hate analysis... The solutions are boring and microscopic.

OpenStudy (blacksteel):

@zzr0ck3r Ah, you are correct. I was just skimming posts and missed the one where you actually wrote the proof. That said, I wrote it the way I did because its easy, without a lot of work, to prove that a rational number will always exist in a neighborhood around 0. (Obviously this can be extrapolated for any neighborhood of Q in R, but I was going for something simple to understand for people who might not have as much set theory background).

OpenStudy (zzr0ck3r):

Note that for any real number there is a sequence of rational numbers that converges to it. So there is a sequence in in our set that converges to the sup of our set. So for any positive number given there is an element of our sequence between our "largest element" and the supremum of our set, and we are done.